OpenStudy (credmond):

Use the addition or subtraction formula to find the exact value of the expression. sin(225)

OpenStudy (mathmate):

Hint: sin(225)=sin(180+45) use the sum formula sin(a+b)=sin(a)cos(b)+cos(a)sin(b). sin(180)=0 sin(45)=\( \frac{\sqrt{2}}{2} \)